A box contains 5 red marbles, 8 white marbles and 4 green marbles. One marble is taken out of the box at random. What is the probability that the marble taken out will be

(i) red?

(ii) white?

(iii) not green?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
The Total no. of balls = 5+8+4 = 17 P(E) = (Number of favourable outcomes/ Total number of outcomes) (i) Total number of red balls = 5 P (red ball) = 5/17 = 0.29 (ii) Total number of white balls = 8 P (white ball) = 8/17 = 0.47 (iii) Total number of green balls = 4 P (green ball) = 4/17 = 0.23 ∴ P (not green) = 1-P(green ball) = 1-(4/7) = 0.77